How many ways can 9 books be arranged on a shelf?
Assuming that each of the nine books is unique, that there are only nine positions open on the given shelf, and that each book can fit in each position, the answer is 9! (nine factorial) which is equal to 362,880.
If the nine books are not all unique (i.e. there are two of the same book), the number should decrease taking into effect that reversing the positions of the identical books makes no overall change.
If there is one additional position, the number should increase to 10! or 3,628,800. If there is more than one additional position, the number will still increase, but not to 11! or 12! since both open spaces are reversible just like the identical books.
What is the probability of drawing a heart a diamond or a spade?
the probibility of drawing a heart diamond or a spaid is a 1/3 chance........
What is the probability of drawing an eight or a spade out of a deck of 52 cards?
Number of cards in a deck =52
Number of cards that are spade =13
Number of cards that are eight =4
Number of cards that are spade and eight =1
Probability that the card drawn is a spade or an eight =
13/52+4/52 -1/52
=16/52 or 4/13
What is the probability of not drawing a diamond in a deck of 52 cards?
number of cards in a deck =52
number of cards that are not diamond =39
Probability that the card drawn is not a diamond= 39/52 = 3/4
What are the steps to find the mean absolute deviation?
add the numbers and divide them by how many numbers there are
What is the probability of not drawing a heart?
Number of cards in a deck = 52
Number of cards that are heart = 13
Therefore number of cards that are not heart = 52-13 = 39
Probability of not drawing a heart = 39/52 or 3/4
How many ways can you choose 2 books from the 6 on the shelf?
20, without replacing the first book chosen.
25 with replacing it.
How many outcomes are possible if a number cube is rolled twice?
The possible outcomes for rolling a number cube twice are:
1,1 1,2 1,3 1,4 1,5 1,6 2,1 2,2 2,3 2,4 2,5 2,6 3,1 3,2 3,3 3,4 3,5 3,6 4,1 4,2 4,3 4,4 4,5 4,6 5,1 5,2 5,3 5,4 5,5 5,6 6,1 6,2 6,3 6,4 6,5 6,6
This is a total of 36 different combinations.
What is the outcome if two coins are tossed and one number cube is rolled?
Let's call one coin A and the other B. omes
The possible outcomes for the coins are; A heads and B tails, A tails and B heads, A and B heads, A and B tails. That's four outcomes.
The possible outcomes for a single die (as in dice) are six since a die has six faces,
So four times six is twenty four possible outcomes.
What is the probability of obtaining exactly 4 tails from the 9 flips of the coin?
It is approx 0.2461
What happens to theoretical and experimental probability when you increase the number of trials?
When you increase the number of trials of an aleatory experiment, the experimental
probability that is based on the number of trials will approach the theoretical
probability.
Is the data is sample or population?
Data is neither sample nor population. Data are collected for attributes. These can be for a sample or a population.
What is the probability of that you roll a multiple of 3 or 5 on a number cube?
The probability that you roll a multiple of 3 (3 and 6) in a fair die is: P(3 or 6) = 2/6
=1/3 = 0.333... ≈ 33.3%.
The probability that you roll a multiple of 5 (5) is: P(5) = 1/6.
The probability that you roll a multiple of 3 or 5 is: P(3 or 6 or 5) = 2/6 + 1/6 = 1/2 = 0.50 = 50%
What is experimental probability of correctly guessing at exactly zero correct answers?
I regret that I do not have access to a study that looked at this matter and so do not have any experimental evidence.
How many permutations are there for the letters in the word telephone?
The number of combinations of the letters TELEPHONE is 9 factorial, or 362,880. From that, however, we have to divide by 4 to account for the E being repeated twice, so the number of distinct combinations is 90,720.
What is the sample space when tossing 3 coins?
The sample space when tossing 3 coins is [HHH, HHT, HTH, HTT, THH, THT, TTH, TTT]
Yes.
How is the term probability related to genetics?
The term probability is related to genetics because they both give guesses about how something that might be the outcome of something
How many possible outcomes would there be if three coins were tossed up three times?
9, you just have to multiply for problems like this
What is the definition of probability in math?
Probability is the likelihood that some particular event will occur. It is expressed as a number between 0 and 1, where 0 means the event will not occur, and 1 means the event will occur. Values in between 0 and 1 indicate the relative likelihood, i.e. "probability", that the event will occur.
For instance, the probability of tossing heads on a fair coin toss is 0.5, the probability of rolling a 3 on a standard die is 1 in 6, or about 0.1667, and the probability of drawing an ace of spades from a standard deck of 52 cards is 1 in 52 or about 0.01923.
How the symmetric distribution is always normal?
The statement is false. The binomial distribution (discrete) or uniform distribution (discrete or continuous) are symmetrical but they are not normal. There are others.